# CoRNStack: High-Quality Contrastive Data for Better Code Ranking

### v2

@@ 1 Introduction
The rapid advancement of software development has led to an increased reliance on automated tools for code generation (Chen et al., 2021; Li et al., 2022b; Nijkamp et al., ). As codebases grow in both size and complexity, the ability to efficiently search for and retrieve relevant code snippets is important. Code retrieval is crucial for advancing Retrieval-Augmented Code Generation (RACG) (Wang et al., 2024b) with large language models (LLMs), where providing contextual examples significantly improves the relevance and accuracy of generated code. Effective code retrieval facilitates bug identification, enables the reuse of existing solutions, and minimizes redundancy, accelerating the overall development process. Code embedding models (Li et al., 2022a; Wang et al., 2023b; Zhang et al., 2024) have gained traction for their ability to encode the semantic and syntactic properties of code into dense vector representations and help retrieve relevant snippets with high precision. However, these approaches have not demonstrated substantial success in real-world applications, particularly in complex tasks like resolving GitHub issues as evaluated by benchmarks like SWE-Bench (Jimenez et al., 2024).
@@ 4.1.2 Results
Table 3 presents the retrieval performance for function-level text-to-code retrieval. Our approach significantly outperforms all open-source and proprietary text and code embedding models, establishing a new state-of-the-art for text-to-code retrieval. Notably, despite being evaluated in a zero-shot setting, our code retriever achieves better performance than CodeT5+, which uses CodeSearchNet for contrastive finetuning. Further, our 137M parameter encoder outperforms the 1.3B CodeSage-Large model, which is ten times larger. Table 4 shows the performance on CoIR, which includes a variety of code retrieval tasks. Our code retriever, despite being smaller than the majority of the baselines, consistently performs well across all the tasks, leading to the highest average performance. This demonstrates the robustness of our contrastive training data, with the trained model exhibiting superior cross-task generalization despite being trained exclusively for only text-to-code retrieval.
Retriever Param. CodeSeachNet AdvTest
Python Java JS PhP Go Ruby Avg. Python
E5-Base 110M 11.2 9.0 8.9 8.6 25.6 16.8 13.4 5.9
Arctic-Embed-M 137M 53.8 49.5 46.3 41.1 71.9 57.9 53.4 34.1
CodeSage-Small 130M 64.4 63.2 60.0 54.7 77.7 63.2 64.9 41.3
CodeSage-Base 356M 68.0 68.0 67.0 58.2 83.2 68.0 68.7 49.1
CodeSage-Large 1.3B 70.8 70.2 69.5 61.3 83.7 71.9 71.2 52.7
Jina-Code-v2 161M 64.4 66.4 61.8 55.9 84.4 70.4 67.2 37.1
CodeT5+ 110M 71.7 71.8 69.2 67.8 90.7 74.4 74.2 40.8
OpenAI-Ada-002 Unknown 68.0 71.5 67.5 60.6 85.6 74.2 71.3 38.1
Voyage-Code-002 Unknown 66.8 64.8 63.4 52.0 88.9 75.0 68.5 -
CodeRankEmbed 137M 78.4 76.9 71.4 68.8 92.7 79.3 77.9 59.5
Table 3: Ranking performance (%) for retrievers of difference sizes on function-level text-to-code retrieval datasets. Following Zhang et al. (2024), we report numbers for MRR@1000.
Retriever Apps CosQA SQL CSN CSN CodeTrans StackOver CodeFeedBack
-CCR -Contest -DL Flow -ST -MT
E5-base 11.5 32.6 52.3 68.0 56.9 62.5 21.9 86.9 74.5 42.0 50.9
Arctic-Embed-M 5.4 27.6 18.9 37.4 62.2 68.9 28 86.9 67 28.0 43.0
E5-Mistral 21.3 51.3 66.0 54.3 65.3 82.6 33.2 91.5 72.7 33.7 55.2
CodeSage-Small 17.3 30.5 51.9 74.1 84.2 76.2 31.0 73.9 62.3 42.6 54.4
CodeSage-Base 27.6 29.4 59.4 76.9 86.9 78.9 31.9 76.2 63.0 44.6 57.5
CodeSage-Large 32.7 28.9 59.5 78.1 89.0 82.6 32.7 78.7 65.4 46.3 59.4
Jina-Code-v2 16.4 42.2 46.4 84.0 82.7 83.6 26.8 89.3 68.6 44.4 58.4
CodeT5+ 3.3 23.1 41.1 78.0 83.6 52.3 31.6 59.9 53.2 32.8 45.9
OpenAI-Ada-002 8.7 29.8 58.3 74.2 69.1 53.3 26.0 72.4 47.1 17.7 45.6
Voyage-Code-002 26.5 29.8 69.3 81.8 73.5 72.8 27.3 77.7 65.4 28.7 56.3
CodeRankEmbed 21.1 36.3 58.8 83.7 86.9 78.8 32.8 82.3 75.7 45.2 60.1
@@ 5 Conclusion
While heuristically filtering examples in CoRNStack, exact matches with downstream evaluation datasets were removed. However, more comprehensive semantic filtering may be necessary to eliminate the possibility of similar queries and code being present. To further validate the functional correctness of the code data, one could generate test cases using a combination of LLMs and/or human annotators. However, considering the scale of the dataset, this approach would be computationally intensive to generate diverse test cases that cover all edge cases and would require substantial resources to set up the necessary environments for execution-based validation. Nevertheless, this is a valuable area for future exploration. Additionally, the benefits of the proposed code retrieval and reranking framework for real-world software development are demonstrated only for function localization, not for downstream code repair performance, due to API costs associated with running GPT-4o on SWE-Bench-Lite.
@@ Acknowledgement
We would like to thank members of the BlenderNLP group and Nomic AI team for valuable feedback and comments. This research is based upon work supported by DARPA ITM Program No. FA8650-23-C-7316 and DARPA SemaFor Program No. HR001120C0123. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of DARPA, or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for governmental purposes notwithstanding any copyright annotation therein.
@@ A.5 Detailed Explanation of Comparison with CodeT5+
Ours 137M Retriever Zero-shot 77.9 59.5

### v3

# CoRNStack: High-Quality Contrastive Data for Better Code Retrieval and Reranking

@@ 1 Introduction
The rapid advancement of software development has led to an increased reliance on automated tools for code generation (Chen et al., 2021; Li et al., 2022b; Nijkamp et al., ; Ugare et al., 2024b; Banerjee et al., 2025; Ugare et al., 2024a; Suresh et al., 2025). As codebases grow in both size and complexity, the ability to efficiently search for and retrieve relevant code snippets is important. Code retrieval is crucial for advancing Retrieval-Augmented Code Generation (RACG) (Wang et al., 2024b) with large language models (LLMs), where providing contextual examples significantly improves the relevance and accuracy of generated code. Effective code retrieval facilitates bug identification, enables the reuse of existing solutions, and minimizes redundancy, accelerating the overall development process. Code embedding models (Li et al., 2022a; Wang et al., 2023b; Zhang et al., 2024) have gained traction for their ability to encode the semantic and syntactic properties of code into dense vector representations and help retrieve relevant snippets with high precision. However, these approaches have not demonstrated substantial success in real-world applications, particularly in complex tasks like resolving GitHub issues as evaluated by benchmarks like SWE-Bench (Jimenez et al., 2024).
@@ 4.1.2 Results
Table 3 presents the code retrieval performance results for CodeSearchNet, AdvTest, and CoIR. The task-specific CoIR results are presented in Table 8 in the Appendix. On CodeSearchNet, our approach significantly outperforms all open-source and proprietary text and code embedding models, establishing a new state-of-the-art for text-to-code retrieval. Notably, despite being evaluated in a zero-shot setting, our code retriever achieves better performance than CodeT5+, which uses CodeSearchNet for contrastive finetuning. Further, our 137M parameter encoder outperforms the 1.3B CodeSage-Large model, which is ten times larger. In addition, on CoIR, our code retriever, despite being smaller than the majority of the baselines, consistently performs well across all the tasks, leading to the highest average performance. This demonstrates the robustness of our contrastive training data, with the trained model exhibiting superior cross-task generalization despite being trained exclusively for only text-to-code retrieval.
Retriever Param. CodeSeachNet AdvTest COIR
Python Java JS PhP Go Ruby Avg. Python Avg.
Arctic-Embed-M 137M 53.8 49.5 46.3 41.1 71.9 57.9 53.4 34.1 43.0
CodeSage-Small 130M 64.4 63.2 60.0 54.7 77.7 63.2 64.9 41.3 54.4
CodeSage-Base 356M 68.0 68.0 67.0 58.2 83.2 68.0 68.7 49.1 57.5
CodeSage-Large 1.3B 70.8 70.2 69.5 61.3 83.7 71.9 71.2 52.7 59.4
Jina-Code-v2 161M 64.4 66.4 61.8 55.9 84.4 70.4 67.2 37.1 58.4
CodeT5+ 110M 71.7 71.8 69.2 67.8 90.7 74.4 74.2 40.8 45.9
OpenAI-Ada-002 Unknown 68.0 71.5 67.5 60.6 85.6 74.2 71.3 38.1 45.6
Voyage-Code-002 Unknown 66.8 64.8 63.4 52.0 88.9 75.0 68.5 - 56.3
CodeRankEmbed 137M 78.4 76.9 71.4 68.8 92.7 79.3 77.9 59.5 60.1
Table 3: Ranking performance (%) for different retrievers on various code retrieval benchmarks. We report the official metrics for each dataset: MRR@1000 for CodeSearchNet and Advtest, and nDCG@10 for COIR. Detailed task-wise numbers for COIR are in Table 8 in Appendix.
@@ 5 Conclusion
Heuristic filtering in CoRNStack removed exact matches with evaluation datasets, but additional semantic filtering is needed to catch similar queries and code. To further validate the functional correctness of the code data, one could generate test cases using LLMs and/or human annotators. However, due to the dataset size, it would be computationally intensive to generate comprehensive test cases and set up the necessary environments for execution-based validation.
@@ Acknowledgement
We would like to thank the BlenderNLP group and Nomic AI team for valuable feedback. This research is based upon work supported by DARPA ITM Program No. FA8650-23-C-7316 and DARPA SemaFor Program No. HR001120C0123. The views and conclusions contained herein are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of DARPA, or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for governmental purposes notwithstanding any copyright annotation therein.
@@ A.2 Accuracy of (text, code) pairings by language
Retriever Apps CosQA SQL CSN CSN CodeTrans StackOver CodeFeedBack
-CCR -Contest -DL Flow -ST -MT
E5-base 11.5 32.6 52.3 68.0 56.9 62.5 21.9 86.9 74.5 42.0 50.9
Arctic-Embed-M 5.4 27.6 18.9 37.4 62.2 68.9 28 86.9 67 28.0 43.0
E5-Mistral 21.3 51.3 66.0 54.3 65.3 82.6 33.2 91.5 72.7 33.7 55.2
CodeSage-Small 17.3 30.5 51.9 74.1 84.2 76.2 31.0 73.9 62.3 42.6 54.4
CodeSage-Base 27.6 29.4 59.4 76.9 86.9 78.9 31.9 76.2 63.0 44.6 57.5
CodeSage-Large 32.7 28.9 59.5 78.1 89.0 82.6 32.7 78.7 65.4 46.3 59.4
Jina-Code-v2 16.4 42.2 46.4 84.0 82.7 83.6 26.8 89.3 68.6 44.4 58.4
CodeT5+ 3.3 23.1 41.1 78.0 83.6 52.3 31.6 59.9 53.2 32.8 45.9
OpenAI-Ada-002 8.7 29.8 58.3 74.2 69.1 53.3 26.0 72.4 47.1 17.7 45.6
Voyage-Code-002 26.5 29.8 69.3 81.8 73.5 72.8 27.3 77.7 65.4 28.7 56.3
CodeRankEmbed 21.1 36.3 58.8 83.7 86.9 78.8 32.8 82.3 75.7 45.2 60.1
@@ A.5 Detailed Explanation of Comparison with CodeT5+
CodeRankEmbed 137M Retriever Zero-shot 77.9 59.5